# Instability of Legendrian knottedness, and non-regular Lagrangian concordances of knots

### v1

@@ 5. Constructing non-regular Lagrangian concordances
First, recall that the space of oriented real two-planes $V\subset\mathbb{R}^{4}$ can be identified with $\mathbb{C}P^{1}\times\mathbb{C}P^{1}\cong Gr^{+}_{2}(\mathbb{R}^{4})$. Here we identify the first $\mathbb{C}P^{1}$-factor with an element in the twistor-space $SO(4)/U(2)$, i.e. an almost complex structure $J\in SO(4)$, while the second factor denotes a $J$-complex line in $\mathbb{R}^{4}$. For an appropriate such identification, the Lagrangian planes are given by the subset
$S^{1}\times\mathbb{C}P^{1}\subset\mathbb{C}P^{1}\times\mathbb{C}P^{1}\cong Gr^{+}_{2}(\mathbb{R}^{4}),$
while the non-totally real 2-planes for the standard complex structure $J_{0}$ can be identified with
$D=\{0,\infty\}\times\mathbb{C}P^{1}\subset\mathbb{C}P^{1}\times\mathbb{C}P^{1}\cong Gr^{+}_{2}(\mathbb{R}^{4}).$
More precisely, the latter submanifold consists of those planes that coincide with $J_{0}$-complex planes equipped either with the complex or anti-complex orientation.
It follows from the above that any smooth concordance $C^{\prime}$ between $\Lambda_{\pm}$, where $\tt{rot}(\Lambda_{\pm})=0$, that is cylindrical outside of a compact subset has a Gauss map of tangent planes
$\Gamma\colon C^{\prime}\to Gr^{+}_{2}(\mathbb{R}^{4})$
that is compactly supported homotopic to a family of totally real planes if and only if the intersection number with the hypersurface $D\subset\mathbb{C}P^{1}\times\mathbb{C}P^{1}\cong Gr^{+}_{2}(\mathbb{R}^{4})$ vanishes. Indeed, the assumptions on the boundary imply that the Gauss map along $\partial C^{\prime}$ maps to a null-homologous curve inside the equator $S^{1}\subset\mathbb{C}P^{1}=SO(4)/SU(2)$ when projected to the first factor in the above identification $\mathbb{C}P^{1}\times\mathbb{C}P^{1}\cong Gr^{+}_{2}(\mathbb{R}^{4})$. Thus, the algebraic number of pre-images of the two different components
$\{0\}\times\mathbb{C}P^{1}\>\sqcup\>\{\infty\}\times\mathbb{C}P^{1}=D$
under the Gauss map can readily be seen to agree. When both of these two numbers vanish (which, under the current assumptions, is the same as the vanishing of the total algebraic number of pre-images of $D$), one can readily find a homotopy of the Gauss map relative boundary that removes all intersections with $D$.
By the $h$-principle for totally real embeddings relative to boundary [20, 27.4.1 and Example 2], we simply need to show that $C^{\prime}$ has a Gauss map that is homotopy relative boundary into the totally real planes in order to ensure the existence of the sought smooth isotopy. Using the argument in the previous paragraph, it thus only remains to show that the reversed cobordism $C^{\prime}$ has a Gauss map for which the count of pre-images of $D$ vanishes algebraically.
Since $C$ is decomposable, this follows from a local computation in the standard model. Indeed, by the definition of decomposability, $C$ can be written as a concatenation of Lagrangian traces of Legendrian isotopies and Lagrangian standard 00 and $1$-handles. The subset of $C^{\prime}$ that corresponds to a Lagrangian trace is clearly smoothly isotopic to a totally real cobordism, since it is even smoothly isotopic to the Lagrangian trace of the inverse isotopy. It thus suffices to compute the intersections with $D$ for a standard Lagrangian 00 and $1$-handle. Here, note that away from a small neighbourhood of the critical point of the standard Lagrangian handle, we can assume that the Lagrangian is cylindrical and thus the reversed cobordism is again Lagrangian. Since $C$ is a cylinder there is an equal number of 00 and $1$-handles. It can be seen that these two types of handles contribute with intersection numbers with $D$ of opposite signs; indeed, this can be checked by putting two Lagrangian 00 and $1$-handles in cancelling position. ∎
@@
Note that totally real concordances define a relation which is obviously reflexive and transitive. From Proposition 5.4 it follows that this relation is not anti-symmetric, and hence totally real concordances do not define a partial order.
@@
By Theorem 5.1 the Lagrangian concordances produced by Corollary 5.6 are never regular. Therefore, from the result of Conway, Etnyre and Tosun [6], it follows that these Lagrangian concordances also are not decomposable in the sense of Chantraine [4].
@@
A decomposable Lagrangian disc filling can be seen as a standard disc filling of a standard Legendrian unknot $U$ concatenated with a decomposable concordance from $U$ to $\Lambda$. We apply Proposition 5.4 produce a totally real concordance from $\Lambda$ to $U$. Theorem 5.3 then produces the sought Lagrangian concordance, by which we first need to add sufficiently many positive and negative stabilisations of to $U$ and $\Lambda$. ∎
@@ Appendix A Classical obstructions to the existence of totally real cobordisms between Legendrian knots
If there is a totally real orientable cobordism $L$ from Legendrian knot $\Lambda_{-}$ to Legendrian knot $\Lambda_{+}$, then
$tb(\Lambda_{+})-tb(\Lambda_{-})=-\chi(\Lambda),$
@@
The proof is completely analogous to the case when the cobordism is Lagrangian. Namely, we perturb the symplectisation coordinate in a compact subset so that it restricts to a Morse function $h\colon S\to\mathbb{R}$ on the cobordism. Then we consider a push-off by $J(\nabla h)$ for an almost complex structure that is cylindrical outside of a compact subset, compatible with the symplectic structure, and for which the cobordism is totally-real. We can assume that $\nabla h=\partial_{t}$ is satisfied outside of a compact subset. It follows that the signed count of intersection points are precisely $-\chi(S)$ (as in the Lagrangian case). The rest follows in the same manner. (Recall that the Thurston-Bennequin invariant is the self-intersection number computed by the push-off of a Legendrian along the vector field $J\partial_{t}$.) ∎

### v2

@@ 5. Constructing non-regular Lagrangian concordances
Outside of a sufficiently large compact subset that contains the non-cylindrical part of $C$, the reversed cobordism $C^{\prime}$ is already Lagrangian, and thus totally real for any compatible almost complex structure. We will analyse the homotopy class of the tangent planes of $C^{\prime}$ inside $\mathbb{R}\times\mathbb{R}^{3}$, in order to show that the map is compactly supported homotopic to one which takes values in Lagrangian tangent planes; the latter are in particular totally real for any compatible almost complex structure. The existence of the smooth isotopy is then a consequence of the $h$-principle for totally real embeddings relative to the boundary [21, 27.4.1 and Example 2].
The assumption of decomposability implies that the concordance consists of $k$ number of standard 0-handles and $k$ number of standard 1-handles of the types described in Section 3.1. We may assume that all 0-handles are attached before the 1-handles. The 0-handles give rise to a number $k$ of small unlinked unknots. Instead of applying Legendrian isotopies to these unknots and $\Lambda_{-}$, and then performing the cusp-connected sums, we are able to fix both $\Lambda_{-}$ and the $k$ unknots, and then attach the 1-handles along long and complicated (e.g. possibly non-trivially linked) attaching arcs; see Subsections 3.1 and 3.3. An example of this procedure is shown in Figure 18.
The decomposable concordance can thus be described by starting with the Legendrian $\Lambda_{-}$ and $k$ unlinked unknots, together with $k$ complicated surgery arcs. Since the cobordism is a connected concordance, it follows that the surgery arcs corresponding to the 1-handles must connect the $k$ unlinked unknots and the negative end $\Lambda_{-}$; the graph whose $k+1$ vertices correspond to the Legendrian knots and $k$ edges corresponding to the surgery arcs is thus a connected tree. After performing a Legendrian regular homotopy (not Legendrian isotopy!) of the surgery arcs, we may assume that the surgery arcs all become unlinked. To see this it suffices to apply the $h$-principle for Legendrian immersions, but one can also see this by an explicit deformation using the front projection. An example of such a regular homotopy is the one that takes the non-trivially linked surgery arc shown in Figure 18 and replaces it with the unlinked surgery arc shown in Figure 19. Consequently, this means that $C$ admits a regular homotopy through Lagrangians that are
  • coinciding with $(-\infty,-A]\times\Lambda_{-}$ in $(-\infty,-A]\times\mathbb{R}^{3}$; and
  • cylindrical (and possibly immersed) in $A,+\infty)\times\mathbb{R}^{3}$,
for $A\gg 0$, to an embedded decomposable Lagrangian concordance $\tilde{C}$ for which the $k$ number of Lagrangian 0-handles can be cancelled with the $k$ number of Lagrangian 1-handles. In other words $\tilde{C}$ is Hamiltonian isotopic to a trivial concordance.
We will only need the following weaker consequence of the above regular homotopy: the Lagrangian gauss map of $C$ admits a compactly supported homotopy to a map that takes values in Lagrangian planes that are everywhere tangent to $\partial_{t}$, i.e. that takes values in cylindrical Lagrangian tangent planes.
Note that the smooth involution $(t,(x,y,z))\mapsto(-t,(x,y,z))$ of the symplectisation $\mathbb{R}\times\mathbb{R}^{3}$ preserves those two-planes that are tangent to $\partial_{t}$, e.g. the cylindrical Lagrangian tangent planes. Since $C$ has a Gauss map that is compactly supported homotopic to a map of cylindrical Lagrangian planes, we conclude that the tangent map of $C^{\prime}$ is also compactly supported homotopic to a map of Lagrangian tangent planes. This shows that the $h$-principle for totally real embeddings can be applied to $C^{\prime}$ as sought. ∎
@@
Note that totally real concordances define a relation on the set of Legendrian isotopy classes of Legendrian knots which is obviously reflexive and transitive. From Proposition 5.4 it follows that this relation is not anti-symmetric, and hence totally real concordances do not define a partial order.
@@
By Theorem 5.1 the Lagrangian concordances produced by Corollary 5.6 are never regular. Therefore, from the result of Conway, Etnyre and Tosun [6] (or, alternatively, from our generalisation Proposition 3.2), it follows that these Lagrangian concordances also are not decomposable in the sense of Chantraine [4].
@@
A decomposable Lagrangian disc filling can be seen as a standard disc filling of a standard Legendrian unknot $U$ concatenated with a concordance from $U$ to $\Lambda$. We apply Proposition 5.4 in order to produce a totally real concordance from $\Lambda$ to $U$. Theorem 5.3 then gives us the sought Lagrangian concordance, by which we first need to add sufficiently many positive and negative stabilisations to $U$ and to $\Lambda$. ∎
@@ Appendix A Classical obstructions to the existence of totally real cobordisms between Legendrian knots
If there is a totally real orientable cobordism $L$ from null-homologous Legendrian knot $\Lambda_{-}$ with a Seifert surface $S$ to Legendrian knot $\Lambda_{+}$, then
$tb(\Lambda_{+})-tb(\Lambda_{-})={\color[rgb]{0,0,0}-\chi(L)},$
@@
The proof is completely analogous to the case when the cobordism is Lagrangian. Namely, we perturb the symplectisation coordinate in a compact subset so that it restricts to a Morse function $h\colon\color[rgb]{0,0,0}L\color[rgb]{0,0,0}\to\mathbb{R}$ on the cobordism. Then we consider a push-off by $J(\nabla h)$ for an almost complex structure that is cylindrical outside of a compact subset, compatible with the symplectic structure, and for which the cobordism is totally-real. We can assume that $\nabla h=\partial_{t}$ is satisfied outside of a compact subset. It follows that the signed count of intersection points are precisely $\chi(\color[rgb]{0,0,0}L)$ (as in the Lagrangian case). The rest follows in the same manner. (Recall that the Thurston-Bennequin invariant is the self-intersection number computed by the push-off of a Legendrian along the vector field $J\partial_{t}$.) ∎